# Conceptualization, Methodology, Software, Analysis, Writing and reviewing manuscript

### v1

@@ 5.2 Differential Growth, Buckling, and Morphogenesis
where the parameters are taken as $k=0.5\,\text{s}^{-1}\cdot\text{kPa}^{-1}$, and $b^{g}_{r}=b^{g}_{\theta}=b^{g}$ in this simulation. This disparity in growth rates can lead to significant residual stresses and mechanical instabilities that shape the tissue morphology over time.
Figure 8: Cylinder growth with different growth ratios in the radial and circumferential directions. (a) Transverse sections over time in the chick foregut, adapted from [52]. (b) Growth morphology at different times with $b^{g}=0.1E_{0}$, (c) $b^{g}=0.2E_{0}$, and (d) $b^{g}=0.3E_{0}$.
Figs. 6b-d illustrate the time evolution of the cross-sectional shapes. We vary the biochemical driving force ($b^{g}$) to represent different levels of differential expansion. For the mild differential growth (Fig. 6b), the cylinder remains largely axisymmetric at early times, but slight radial undulations emerge as growth progresses. The folds are shallow, indicating that while differential growth initiates instability, the magnitude of the growth mismatch is not sufficient to induce large-amplitude wrinkling. The moderate differential growth (Fig. 6c) shows more pronounced folding patterns. The radial expansion outpaces the circumferential expansion, creating residual hoop stresses that drive the tissue to buckle inward and outward periodically around the circumference. In Fig. 6d, the shell rapidly develops deep, well-defined folds at relatively early times ($t\simeq\tau$). During the differential growth, the number of folds may also increase. This bifurcation is common in tissue development [41, 10]. Fig. 6a depicts cross-sectional slices of a chick foregut at different embryonic days. The progressive formation of circumferential folds resembles the morphologies observed in our PINN simulations under moderate to high differential growth. Although the exact geometry and material properties of embryonic tissues are far more complex, often involving multilayered structures and active cellular processes, the PINN-based model provides a mechanistic view of how differential growth can induce wrinkling and folding patterns reminiscent of in vivo development.
@@ 6 Discussion
A key highlight of our work is the natural emergence of buckling instabilities in creep simulations. Unlike traditional methods that typically rely on artificially imposed imperfections to trigger buckling, the inherent training dynamics of the neural network optimizer exploring the energy landscape appear to serve as a built-in mechanism for capturing the transition from stable creep to buckled configurations. This characteristic is particularly advantageous when modeling viscoelastic structures, where the gradual accumulation of strain leads to complex instability patterns. Furthermore, the extension of our framework to simulate tissue growth and morphogenesis underscores its versatility. By incorporating differential growth rates, the model successfully replicates the formation of intricate folding patterns reminiscent of those observed in biological tissues, thereby providing valuable insights into the mechanics of morphogenesis and offering potential applications in tissue engineering and bio-inspired design.
Despite these promising advances, the study also reveals several limitations that warrant further investigation. For instance, while the framework effectively captures a buckling mode through simulation, it is currently not equipped to conduct a comprehensive eigenvalue buckling analysis as performed by traditional finite element methods, which can identify multiple potential buckling modes and their associated critical conditions. The natural perturbation is a strong, evidence-based hypothesis that needs further theoretical investigation into the link between optimizer dynamics and bifurcation theory. Additionally, the iterative nature of the incremental viscoelastic training loop, particularly when fine time-step resolution is required, may lead to significant computational cost compared to highly optimized implicit FEM solvers. These suggest that further optimization of the training process—possibly through adaptive time-stepping strategies or enhanced neural network architectures—will be necessary to fully exploit the potential of PINN in highly nonlinear or multiscale scenarios.
@@ Acknowledgments
Support from the National Natural Science Foundation of China (Grant nos. 12032014 and T2488101) is acknowledged.
@@ Appendix A Validation Benchmark: Buckling Mode of a Hyperelastic Cantilever Beam
To verify the fundamental capability of the implemented energy-based PINN framework to capture structural buckling, we simulated the classic Euler buckling of a hyperelastic cantilever beam under axial compression. The cantilever beam geometry (length $L=10$ m, height $h=1$ m, assuming plane strain) and boundary conditions (clamped at $x=0$, axial pressure applied at $x=L$) were identical to those described in Section 4.1 of the main text. The material was modeled as compressible Neo-Hookean hyperelastic (described by Eqs. (1) and (2)), using the instantaneous elastic properties of the viscoelastic material: Young’s Modulus $E=E_{1}+E_{\infty}=10$ Pa and Poisson’s ratio $\nu=0.35$. The theoretical Euler critical buckling pressure for this configuration can be calculated by Eq. (13) , which is $2.06\times 10^{-2}$ Pa. The eigenvalue analysis of FEM (with ABAQUS, Buckle step, element CPE8R) is $2.34\times 10^{-2}$ Pa for first buckling mode and $0.203$ Pa for second buckling mode.
@@ Appendix B Simulation under Sub-Critical Pressure
To further investigate the role of neural network training dynamics in capturing buckling phenomena, we performed an additional simulation of the hyperelastic cantilever beam under a load known to be below the critical buckling threshold for the simulated time frame. The geometry, material properties, boundary conditions, and the energy-based incremental PINN framework were identical to those used for the buckling analysis in Appendix A. The applied compressive pressure is set to a sub-critical value of $p=1\times 10^{-2}$ Pa, which is smaller than the critical buckling pressure ($p_{cr}=2.34\times 10^{-2}$ Pa).
Figure S2: The deformed cantilever beam when the compressive pressure is lower than the buckling threshold. (a) The displacement in the y-direction of the deformed beam. (b) The convergence plot of the training process.
Fig. S2a shows the predicted displacement in the $y-$direction of the cantilever beam. Unlike the simulation under the critical load (Fig. 5c), there is no sign of rapid divergence or acceleration in deflection characteristic of buckling instability. The beam only develops some oscillatory deformation under perturbation. Fig. S2b shows a representative convergence plot (the loss vs. training iterations). Oscillations during the optimization process are still evident, similar in nature to those observed during the buckling simulations.

### v2

@@ 5.2 Differential Growth, Buckling, and Morphogenesis
where the parameters are taken as $k=0.5\,\text{s}^{-1}\cdot\text{kPa}^{-1}$, and $b^{g}_{r}=b^{g}_{\theta}=b^{g}$ in this simulation. This disparity in growth rates can lead to significant residual stresses and mechanical instabilities that shape the tissue morphology over time. During the simulation, the stress magnitude is set no greater than $b^{g}$.
Figure 8: Cylinder growth with different growth ratios at the radial and circumferential directions. (a) Transverse sections over time in the chick foregut, adapted from [Gill.2024]. (b) Growth morphology at different times with $b^{g}=0.1E_{0}$, (c) $b^{g}=0.2E_{0}$, and (d) $b^{g}=0.3E_{0}$.
Figs. 6b-d illustrate the time evolution of the cross-sectional shapes. We vary the biochemical driving force ($b^{g}$) to represent different levels of differential expansion. For the mild differential growth (Fig. 6b), the cylinder remains largely axisymmetric at early times, but slight radial undulations emerge as growth progresses. The folds are shallow, indicating that while differential growth initiates instability, the magnitude of the growth mismatch is not sufficient to induce large-amplitude wrinkling. The moderate differential growth (Fig. 6c) shows more pronounced folding patterns. The radial expansion outpaces the circumferential expansion, creating residual hoop stresses that drive the tissue to buckle inward and outward periodically around the circumference. In Fig. 6d, the shell rapidly develops deep, well-defined folds at relatively early times ($t\simeq\tau$). During the differential growth, the number of folds may also increase. This bifurcation is common in tissue development [Li.2011, BenAmar.2025]. Fig. 6a depicts cross-sectional slices of a chick foregut at different embryonic days. The progressive formation of circumferential folds resembles the morphologies observed in our PINN simulations under moderate to high differential growth. Although the exact geometry and material properties of embryonic tissues are far more complex, often involving multilayered structures and active cellular processes, the PINN-based model provides a mechanistic view of how differential growth can induce wrinkling and folding patterns reminiscent of in vivo development.
@@ 6 Discussion
A key highlight of our work is the natural emergence of buckling instabilities in creep simulations. Unlike traditional methods that typically rely on artificially imposed imperfections to trigger buckling, the inherent training dynamics of the neural network optimizer exploring the energy landscape appear to serve as a built-in mechanism for capturing the transition from stable creep to buckled configurations. This characteristic is particularly advantageous when modeling viscoelastic structures, where the gradual accumulation of strain leads to complex instability patterns. Furthermore, the extension of our framework to simulate tissue growth and morphogenesis underscores its versatility. By incorporating differential growth rates, the model successfully replicates the formation of intricate folding patterns reminiscent of those observed in biological tissues, thereby providing valuable insights into the mechanics of morphogenesis and offering potential applications in tissue engineering and bio-inspired design. In this study, we also using Fourier feature embeddings into the PINN. The results show that the FF-PINN does not significantly improve the accuracy or convergence for the viscoelastic creep-buckling problems considered in this work. We think the main reason is that the dominant physical behavior involves low-order global buckling modes mainly, which do not require high-frequency representational capacity.
Despite these promising advances, the study also reveals several limitations that warrant further investigation. For instance, while the framework effectively captures a buckling mode through simulation, it is currently not equipped to conduct a comprehensive eigenvalue buckling analysis as performed by traditional finite element methods, which can identify multiple potential buckling modes and their associated critical conditions. The natural perturbation is a strong, evidence-based hypothesis that needs further theoretical investigation into the link between optimizer dynamics and bifurcation theory. Additionally, the iterative nature of the incremental viscoelastic training loop, particularly when fine time-step resolution is required, may lead to significant computational cost compared to highly optimized implicit FEM solvers. These suggest that further optimization of the training process—possibly through adaptive time-stepping strategies or enhanced neural network architectures—will be necessary to fully exploit the potential of PINN in highly nonlinear or multiscale scenarios. Moreover, operator learning would amortize the computational cost of many-query tasks (design optimization, uncertainty quantification, digital twins), but applying it to nonlinear, history-dependent viscoelasticity introduces new challenges: the input space must represent the past deformation history or internal variables, training requires large, representative ensembles of solutions across parameter and history space, and careful incorporation of energy/thermodynamic constraints is necessary to preserve physical consistency. In future work, we can investigate strategies to represent memory in operator architectures, to reduce data requirements via transfer learning, and to combine energy-based losses with operator training to retain stability and thermodynamic consistency.
@@ Acknowledgments
Support from the National Natural Science Foundation of China (Grant nos. 12032014, T2488101 and 12502234 ) is acknowledged.
@@ Appendix A Validation Benchmark: Buckling Mode of a Hyperelastic Cantilever Beam
To verify the fundamental capability of the implemented energy-based PINN framework to capture structural buckling, we simulated the classic Euler buckling of a hyperelastic cantilever beam under axial compression. The cantilever beam geometry (length $L=10$ m, height $h=1$ m, assuming plane strain) and boundary conditions (clamped at $x=0$, axial pressure applied at $x=L$) were identical to those described in Section 4.1 of the main text. The material was modeled as compressible Neo-Hookean hyperelastic (described by Eqs. (1) and (2)), using the instantaneous elastic properties of the viscoelastic material: Young’s Modulus $E=E_{1}+E_{\infty}=10$ Pa and Poisson’s ratio $\nu=0.35$. The theoretical Euler critical buckling pressure for this configuration can be calculated by Eq. (13) , which is $2.06\text{e}-2$ Pa. The eigenvalue analysis of FEM (with ABAQUS, Buckle step, element CPE8R) is $2.34\text{e}-2$ Pa for first buckling mode and $0.203$ Pa for second buckling mode.
@@ Appendix B Simulation under Sub-Critical Pressure
To further investigate the role of neural network training dynamics in capturing buckling phenomena, we performed an additional simulation of the hyperelastic cantilever beam under a load known to be below the critical buckling threshold for the simulated time frame. The geometry, material properties, boundary conditions, and the energy-based incremental PINN framework were identical to those used for the buckling analysis in Appendix A. The applied compressive pressure is set to a sub-critical value of $p=1\text{e}-2$ Pa, which is smaller than the critical buckling pressure ($p_{cr}=2.34\text{e}-2$ Pa).
Figure S2: The deformed cantilever beam when the compressive pressure is lower than the buckling threshold. (a) The displacement in the y-direction of the deformed beam. (b) The convergence plot of training process.
Fig. S2a shows the predicted displacement at the $y$-direction of the cantilever beam. Unlike the simulation under the critical load (Fig. 5c), there is no sign of rapid divergence or acceleration in deflection characteristic of buckling instability. The beam only develops some oscillatory deformation under perturbation. Fig. S2b shows a representative convergence plot (the loss vs. training iterations). Oscillations during the optimization process are still evident, similar in nature to those observed during the buckling simulations.
@@
Appendix C Natural dynamics of Adam optimizer
The Adam optimizer combines the momentum algorithm and adaptive scaling scheme during training [kingma2014adam]. It updates the parameters as follows:
$\theta_{t+1}=\theta_{t}-\frac{\alpha}{\sqrt{\hat{v_{t}}}+\epsilon}\cdot\hat{m_{t}}$, (31)
where $\theta$ represents the trainable parameters in neural networks, $\alpha$ represents the learning rate, $\epsilon$ is a numerical stabilizer that prevents division by zero and is normally of $1\times 10^{-8}$, $\hat{m_{t}}$ and $\hat{v_{t}}$ are the first and second momentum, respectively. The two momentum factors are updated by
$\hat{m_{t}}$ $=\frac{m_{t}}{1-\beta_{1}^{t}},$ (32)
$\hat{v_{t}}$ $=\frac{v_{t}}{1-\beta_{2}^{t}},$
where
$m_{t+1}$ $=\beta_{1}\cdot m_{t}+(1-\beta_{1})\cdot g_{t},$ (33)
$v_{t+1}$ $=\beta_{2}\cdot v_{t}+(1-\beta_{2})\cdot g_{t}^{2}.$
In Eq.33, $\beta_{1}$ and $\beta_{2}$ are two parameters in the Adam optimisers and are defaultly set as 0.9 and 0.999, $g_{t}$ is the gradient of loss function, e.g., the overall energy of the system
$g_{t}=\nabla_{\theta}\Pi.$ (34)
It is clear to find that the correction of the two momenta is effective at the beginning of training. Given that the training of PINN for viscoelastic problems require large number of training epochs, we simplify and ignore Eq.32 in the following analysis.
Since $m_{t}$ and $v_{t}$ are all initialised as zero vectors, Eq.33 can be also expressed as
$m_{t+1}$ $=(1-\beta_{1})\sum_{n=0}^{t}{\beta_{1}}^{n}\cdot g_{t-n}$ (35)
$=(1-\beta_{1})\sum_{n=0}^{t}{\beta_{1}}^{n}\cdot\nabla_{\theta}\Pi_{t-n}$
$v_{t+1}$ $=(1-\beta_{2})\sum_{n=0}^{t}{\beta_{2}}^{n}\cdot g_{t-n}^{2}$
$=(1-\beta_{2})\sum_{n=0}^{t}{\beta_{2}}^{n}\cdot(\nabla_{\theta}\Pi_{t-n})^{2}.$
By substituting Eq.35 into Eq.31, one can obtain
$\theta_{t+1}$ $=\theta_{t}-\frac{\alpha}{\sqrt{v_{t}}+\epsilon}\cdot m_{t+1}$ (36)
$=-\frac{\alpha}{\sqrt{v_{t}}+\epsilon}\cdot(1-\beta_{1})\sum_{n=0}^{t}{\beta_{1}}^{n}\cdot(\nabla_{\theta}\Pi_{t-n})$
$=-\frac{\alpha(1-\beta_{1})}{\sqrt{v_{t}}+\epsilon}\cdot(\nabla_{\theta}\Pi_{t}+\sum_{n=1}^{t}{\beta_{1}}^{n}\cdot\nabla_{\theta}\Pi_{t-n})$
Let $\tau_{t}=\frac{\sqrt{v_{t}}+\epsilon}{(1-\beta_{1})}$ and $\Phi_{t}=-\sum_{n=1}^{t}{\beta_{1}}^{n}\cdot\nabla_{\theta}\Pi_{t-n}$, Eq.36 can be simplifed and formulated as
$\theta_{t+1}-\theta_{t}=-\frac{\alpha}{\tau_{t}}\cdot(\nabla_{\theta}\Pi_{t}-\Phi_{t}).$ (37)
Eq.37 is the stepwise parameter update rule. The continuous version of Eq.37 (when $\alpha\to 0$) [wang2021understanding] is given as
$\tau_{t}\cdot\dot{\theta}_{t}=-\nabla_{\theta}\Pi_{t}+\Phi_{t},$ (38)
$\Rightarrow$ $0=-\nabla_{\theta}\Pi_{t}-\tau_{t}\cdot\dot{\theta}_{t}+\Phi_{t}.$
This equation is similar to the governing equation of the Langevin dynamics, which reads
$M\ddot{x}=-\nabla U(x)-\gamma\dot{x}+R_{t},$ (39)
where $M$ is the mass of a particle in the Langevin dynamics, $U(x)$ denotes a potential, $R_{t}$ is a random white noise satisfying the Gaussian distribution. As observed, Eq.38 lacks the inertia term compared to Eq.39 and therefore is known as the underdamped condition. During the training processes, the trainable parameters are updated by the Langevin dynamics, approching the minimal of energy potential. Meanwhile, the weighted mean of previous energy gradients, $\Phi_{t}$, acts as a perturbation during the training, preventing the trainable paramters trapped by saddle points. Moreover, $\tau$ is equivalent to the damper $\gamma$. From the dynamics point of view, this damper stabilises training and reduces loss oscillations. From the viewpoint of deep learning, $\tau$ is also an adaptive scaling vectors that balance the learning rate of different trainable parameters, assigning slowly-changing parameters a small damper and a large damper for frequently-changing parameters.